OpenStudy (anonymous):

2/3 +1/4 I need to calculate the sum? I would think the sum is 3/12. This is incorrect?

OpenStudy (unklerhaukus):

\[\frac 23 +\frac14=\frac{2\times4}{3\times4}+\frac{1\times3}{4\times3}=\frac{2\times4+1\times3}{12}\]

OpenStudy (precal):